P10.6 :

Please explain/clarify variable meanings too!
Answer #1

Expand the commutator on left hand side of the equation.

[\hat{S_{x}},\hat{S_{y}}] = \hat{S_{x}}\hat{S_{y}} - \hat{S_{y}}\hat{S_{x}}

State \alpha is the spin up eigen state of \hat{S_{z}} , the z component of the Spin operator. \hat{S_{x}} and \hat{S_{y}} are the x and y components respectively of the spin operators, all of which acts on the spin up and spin down states \left ( \beta \right ) as given in P10.6 except for \hat{S_{y}} acting on \beta . The correct expression is given below. Note this correction.

\hat{S_{y}}\beta = - i$\hbar$\alpha /2  

Act the commutator on the state \alpha

[\hat{S}_{x},\hat{S}_{y}] \alpha = \hat{S}_{x}(\hat{S}_{y}\alpha) - \hat{S}_{y}(\hat{S}_{x}\alpha)

\Rightarrow [\hat{S}_{x},\hat{S}_{y}] \alpha =\hat{S}_{x}(i$\hbar$\beta)/2 - \hat{S}_{y}($\hbar$\beta)/2

\Rightarrow [\hat{S}_{x},\hat{S}_{y}] \alpha =i$\hbar$(\hat{S}_{x}\beta)/2 - $\hbar$ (\hat{S}_{y}\beta)/2

\Rightarrow [\hat{S}_{x},\hat{S}_{y}] \alpha =i$\hbar$($\hbar$ \alpha/2)/2 - $\hbar$(-i$\hbar$\alpha/2)/2

\Rightarrow [\hat{S}_{x},\hat{S}_{y}] \alpha =i$\hbar$^{2} \alpha/4 + i$\hbar$^{2}\alpha/4 = i$\hbar$^{2}\alpha/2

Now note the action of \hat{S_{z}} on \alpha

$\hbar$\alpha /2 = \hat{S_{z}}\alpha

Substitute this result in the evaluated answer for the commutator and the desired result is obtained

[\hat{S}_{x},\hat{S}_{y}] \alpha = i$\hbar$\hat{S}_{z}